# Long-Context Generalization with Sparse Attention

### v1

@@ 1 Introduction
In this paper, we address the root cause of these problems by replacing softmax with $\alpha$-entmax (Peters et al., 2019), a differentiable sparse transformation that induces probability distributions where irrelevant tokens receive exactly zero attention. While $\alpha$-entmax has been used successfully in transformers (Correia et al., 2019), its length generalization properties, to the best of our knowledge, have never been studied. We show theoretically and empirically that $\alpha$-entmax consistently helps to address challenges in long context modeling. Our key contributions include: [^1] Our code is available at https://github.com/deep-spin/bigfish.
@@ 2.1 Transformers
where $p_{ij}^{(\ell)}$ are attention weights, $\text{FFN}^{(\ell)}$ is the feed-forward network, $\text{norm}(\cdot)$ represent LayerNorm modules (Xiong et al., 2020). The output is computed as $\bm{y}_{i}=\text{norm}_{3}\left(\bm{v}^{(L)}_{i}\right)$. The attention weights $p_{ij}^{(\ell)}=\pi(\bm{z}_{i}^{(\ell)})_{j}$ are computed by applying a transformation $\pi:\mathbb{R}^{n}\to\triangle_{n}$ to the attention logits $z_{ij}^{(\ell)}=\langle\bm{q}_{i}^{(\ell)},\bm{k}_{j}^{(\ell)}\rangle/\sqrt{d}$, where $\triangle_{n}=\{\bm{p}\in\mathbb{R}^{n}\,:\,\sum_{i=1}^{n}p_{i}=1,\,\bm{p}\geq\mathbf{0}\}$ represents the probability simplex. Standard transformers employ the softmax function as $\pi$. In this work, we study transformers by casting $\pi$ as the $\alpha$-entmax transformation.
@@ 2.2 $\alpha$-entmax
where $[\cdot]_{+}=\max(0,\cdot)$ and $\tau:\mathbb{R}^{n}\to\mathbb{R}$ yields a threshold that ensures the resulting distribution sums to $1$. A key property of $\alpha$-entmax is that tokens with scores below the threshold receive exactly zero probability, creating sparse attention patterns. When $\alpha\rightarrow 1^{+}$, this reduces to the standard softmax function. The sparsity level increases with $\alpha$, with $\alpha=2$ corresponding to the sparsemax function (Martins & Astudillo, 2016). Figure 2 illustrates $\alpha$-entmax for different values of $\alpha$. We provide more information on $\alpha$-entmax in §A. While $\alpha$-entmax is a suitable choice for sparse attention, its theoretical and empirical impact on long inputs is still unclear. In the next section, we demonstrate how it fundamentally changes the way attention behaves for long contexts.
@@ 3.2 Attention Dispersion and Concentration
These definitions allow us to precisely characterize how softmax and $\alpha$-entmax behave as sequence length increases:
@@
The full proof can be consulted in §D. The key takeaway of this theorem is that the entropy of attention distributions reveals how concentrated or dispersed they are across tokens. While softmax distributions approach maximum entropy $\Theta(\log n)$ as the sequence length increases (indicating complete dispersion), $\alpha$-entmax distributions maintain bounded entropy $\mathcal{O}(\log s)$ where $s$ is the support size. This allows models with $\alpha$-entmax to maintain focused, low-entropy attention patterns even when processing extremely long sequences, as long as the support size is smaller than the full sequence length. This non-dispersion property means transformers with $\alpha$-entmax can scale to very long contexts without the attention becoming dispersed, maintaining their ability to focus on relevant information regardless of how much additional context is present.
@@ 3.3 Representational Preservation and Over-squashing Alleviation
Two critical challenges in long-context transformers are representational collapse and over-squashing—both fundamentally linked to the properties of softmax attention (Barbero et al., 2024).
@@
[Representational Collapse and Over-squashing] A softmax-based transformer suffers from representational collapse when token representations become similar as sequence length grows: for any $\epsilon>0$, there exists $n_{\epsilon}$ such that $\|\bm{v}_{i}^{(L)}-\bm{v}_{j}^{(L)}\|_{1}<\epsilon$ for all $i,j\leq n$ when $n>n_{\epsilon}$.
@@
We show in §C.2 and §C.3 that $\alpha$-entmax addresses both issues: {proposition}[Representational Preservation] There are choices of $\bm{v}^{(0)}\in\mathbb{R}^{n\times d}$ and $\bm{v}^{*(0)}\in\mathbb{R}^{(n+1)\times d}$ such that $\alpha$-entmax ($\alpha>1$) can maintain distinct token representations even as $n\to\infty$: $\|\bm{v}_{n}^{(L)}-\bm{v}_{n+1}^{*(L)}\|_{1}\geq c$ for any constant $c>0$.
@@
[Over-squashing Alleviation] With support size $s\ll n$, $\alpha$-entmax ($\alpha>1$) reduces gradient paths from $\mathcal{O}(n^{L})$ to $\mathcal{O}(s^{L})$, alleviating over-squashing via stronger gradient signals.
To empirically validate these properties, we conducted controlled experiments with transformer architectures employing different attention mechanisms, fully described in §C. For representational collapse, we measured how token representations change when sequence length increases. For over-squashing, we analyzed gradient flow through an 8-layer network on a delayed copying task, where the model must copy information across long distances. The results are shown in Figure 3. The left panel shows the $L_{1}$ norm of representation differences between original and extended sequences after multiple transformer layers. With softmax, this difference rapidly approaches zero as sequence length increases. In contrast, $\alpha$-entmax maintains bounded differences even at extreme lengths (128K tokens), with higher $\alpha$ values providing stronger preservation. The right panel illustrates gradient flow strength across sequence lengths, showing how $\alpha$-entmax maintains significantly stronger gradient signals compared to softmax.
@@ 4.1 Controlling Sparsity in Long Contexts via ASEntmax
To address the challenges discussed above, we propose Adaptive-Scalable $\alpha$-entmax (ASEntmax), which adaptively scales the temperature for each attention head and query based on the context size and content:
@@
where $\bm{w}_{\beta},\bm{w}_{\gamma}\in\mathbb{R}^{d}$ are learnable, head-specific projection vectors. The scaling factor $(\log n)^{\gamma}$ in (9) is directly motivated by the asymptotic behavior of extreme values in normal distributions from Remark 4.1. This allows the model learn a slowly rising ($\gamma>0$) or dampening ($\gamma<0$) temperature schedule without interfering in the positional encodings (which would happen with negative values of $\beta$). Specifically, for IID Gaussian logits $\mathcal{N}(0,\sigma)$, when $\delta=0$ and $\gamma=-0.5$ the scaling counteracts the growth of logit ranges ($\Delta_{n}$) as context increases:
@@
which remains constant as $n$ increases, preventing excessive sparsification. Moreover, ASEntmax recovers standard $\alpha$-entmax when $\beta=0$, hence allowing a smooth transition between scaled and unscaled regimes as context length changes. By making $\bm{w}_{\beta}$ and $\bm{w}_{\gamma}$ head-specific and learnable, ASEntmax can adapt to the optimal scaling behavior for each head, balancing the natural concentration benefits of $\alpha$-entmax with precise control over how sparsity patterns evolve with sequence length. Furthermore, simply scaling the query-key products is appealing from a practical perspective since it allows the direct use of fast optimized kernels for $\alpha$-entmax, such as AdaSplash (Gonçalves et al., 2025), without any modifications.
@@ 4.2 Interacting with Positional Encoding
Positional bias is known to affect attention dynamics and generalization (Barbero et al., 2025; Wu et al., 2025). In this work, we also introduce NAPE (NoPE + ALiBi Positional Encoding), a hybrid scheme that splits attention heads into two groups: some receive no positional encoding (NoPE), allowing flexible, data-driven position inference; others use a fixed monotonic bias (ALiBi) to promote locality. This complementary design allows some heads to focus on global patterns while others encode recency, combining the strengths of both.
Our analysis in §H.1 confirms that models with NoPE can indeed develop relative positional encoding, supporting the flexibility hypothesis of Kazemnejad et al. (2023). While NAPE is not central to our theoretical results, we find it plays a crucial empirical role in enabling robust length generalization, especially when combined with the sparse inductive bias of $\alpha$-attention. Notably, our analysis shows that $\alpha$-entmax interacts with positional encodings in ways that fundamentally alter attention behavior compared to softmax (e.g., creating hard attention windows with ALiBi or inducing frequency-dependent patterns with RoPE). A comprehensive theoretical and empirical comparison between positional encoding methods is provided in §E.
@@ Discussion.
Task-specific patterns further validate our theoretical claims. The consistent superiority of ASEntmax over basic $\alpha$-ntmax confirms the benefits of adaptive scaling, particularly at extreme lengths where fixed-$\alpha$ may become too sparse or too diffuse. SSMax performs well on the Copy task, even outperforming other methods, but struggles on more complex tasks like MQMTAR and Reverse at extreme lengths. This indicates that while scaling logits helps maintain peak attention magnitude, the explicit sparsity of $\alpha$-entmax provides additional benefits by completely removing irrelevant connections. These findings are further supported by results on the Max Retrieval task (Figure 5), where sparse attention mechanisms demonstrate superior length extrapolation compared to dense approaches, with ASEntmax maintaining over 60% accuracy even at 4096-length sequences—a dramatic improvement over standard Softmax and Adaptive Temperature (Veličković et al., 2024).
Finally, Copy and Reverse show moderate generalization (up to 64$\times$ and 8$\times$ respectively), while Sort fails beyond 4$\times$ length for all methods. This pattern suggests that tasks requiring precise global ordering (Sort, Reverse) are inherently more challenging for length generalization than tasks dependent on local or independent token properties. We provide per-task results, including for other positional encoding methods such as ALiBi and NoPE, in §H.
@@ 6 Related Works
The interaction between attention mechanisms and positional encodings significantly impacts long-context generalization. Wu et al. (2025) show that causal masking with distance bias creates a position advantage for early tokens, while Barbero et al. (2025) analyze frequency components in RoPE (Su et al., 2024). Our work shows how $\alpha$-entmax transforms these biases: with NoPE, it helps attention to not always converge to early tokens; with ALiBi, it creates hard attention windows—similar to Hard-ALiBi (Jelassi et al., 2024); with RoPE, it induces frequency-dependent sparsity. Based on this analysis, we introduce NAPE (NoPE + ALiBi), which combines ALiBi’s consistent recency bias with NoPE’s flexible content-based attention for improved long-context generalization.
@@
Previous sparse attention approaches include structured patterns like Longformer (Beltagy et al., 2020), BigBird (Zaheer et al., 2020), Flash-Attention (Dao et al., 2022), as well as adaptive methods like $\alpha$-entmax (Peters et al., 2019), recently accelerated by AdaSplash (Gonçalves et al., 2025). Our work differs by providing a theoretical analysis of why sparsity helps with long-context generalization, linking it to fundamental limitations including dispersion, representational collapse, and over-squashing.
@@ 7 Conclusions
In this paper, we present a principled approach to long-context modeling by replacing softmax with $\alpha$-entmax in transformer attention. Our theoretical analysis demonstrates how this simple change addresses three fundamental limitations: it prevents representational collapse by maintaining distinct token representations, alleviates over-squashing by reducing gradient paths from $\mathcal{O}(n^{L})$ to $\mathcal{O}(s^{L})$, where $s\ll n$ is the number of tokens with nonzero probability, and avoids attention dispersion through naturally sparse distributions. We further introduce Adaptive-Scalable $\alpha$-entmax (ASEntmax), which adaptively adjusts sparsity based on sequence length for each attention head and query input. Our empirical results confirm these theoretical predictions, with ASEntmax significantly outperforming softmax and existing alternatives across diverse tasks, maintaining high performance even at sequence lengths 1000$\times$ beyond training examples. These findings suggest that addressing the fundamental mathematical limitations of transformer attention mechanisms may provide a direct path to achieve robust long-context generalization.
While our experiments focus on controlled settings with small transformers, the theoretical principles we establish should extend to deeper architectures and larger scales; however, future research on sparse attention in production-scale models is still needed, especially with multi-phase approaches commonly employed for extending the context-length in traditional LLMs (Grattafiori et al., 2024).

### v2

@@ 1 Introduction
In this paper, we address the root cause of these problems by replacing softmax with $\alpha$-entmax (Peters et al., 2019), a differentiable sparse transformation that induces probability distributions where irrelevant tokens receive exactly zero attention. While $\alpha$-entmax has been used successfully in transformers (Correia et al., 2019), its length generalization properties, to the best of our knowledge, have never been studied. We show theoretically and empirically that $\alpha$-entmax consistently helps to address challenges in long context modeling. Our key contributions include: [^1] Our code is available at https://github.com/deep-spin/adasplash.
@@ 2.1 Transformers
where $p_{ij}^{(\ell)}$ are attention weights, $\text{FFN}^{(\ell)}$ is the feed-forward network, $\text{norm}(\cdot)$ represent LayerNorm modules (Xiong et al., 2020). The output is computed as $\bm{y}_{i}=\text{norm}_{3}\left(\bm{v}^{(L)}_{i}\right)$. The attention weights $p_{ij}^{(\ell)}=\pi(\bm{z}_{i}^{(\ell)})_{j}$ are computed by applying a transformation $\pi:\mathbb{R}^{n}\to\triangle_{n}$ to the attention logits $z_{ij}^{(\ell)}=\langle\bm{q}_{i}^{(\ell)},\bm{k}_{j}^{(\ell)}\rangle/\sqrt{d}$, where $\triangle_{n}:=\{\bm{p}\in\mathbb{R}^{n}\,:\,\bm{p}\geq\mathbf{0},\mathbf{1}^{\top}\bm{p}=1\}$ represents the probability simplex. Standard transformers employ the softmax function as $\pi$. In this work, we study transformers by casting $\pi$ as the $\alpha$-entmax transformation.
@@ 2.2 $\alpha$-entmax
where $[\cdot]_{+}:=\max(0,\cdot)$ and $\tau:\mathbb{R}^{n}\to\mathbb{R}$ yields a threshold that ensures the resulting distribution sums to $1$. A key property of $\alpha$-entmax is that tokens with scores below the threshold receive exactly zero probability, creating sparse attention patterns. When $\alpha\rightarrow 1^{+}$, this reduces to the standard softmax function. The sparsity level increases with $\alpha$, with $\alpha=2$ corresponding to the sparsemax function (Martins & Astudillo, 2016). Figure 2 illustrates $\alpha$-entmax for different values of $\alpha$. We provide more information on $\alpha$-entmax in §A. While $\alpha$-entmax is a suitable choice for sparse attention, its theoretical and empirical impact on long inputs is still unclear. In the next section, we demonstrate how it fundamentally changes the way attention behaves for long contexts.
@@ 3.2 Attention Dispersion and Concentration
These definitions allow us to examine how softmax and $\alpha$-entmax behave as sequence length grows:
@@
The full proof can be consulted in §D. The key takeaway of this result is that the entropy of attention distributions reveals how concentrated or dispersed they are across tokens. While softmax distributions approach maximum entropy $\Theta(\log n)$ as the sequence length increases (indicating complete dispersion), $\alpha$-entmax distributions maintain bounded entropy $\mathcal{O}(\log s)$ where $s$ is the support size. This allows models with $\alpha$-entmax to maintain focused, low-entropy attention patterns even when processing extremely long sequences, as long as the support size is smaller than the full sequence length. This non-dispersion property means transformers with $\alpha$-entmax can scale to very long contexts without the attention becoming dispersed, maintaining their ability to focus on relevant information regardless of how much additional context is present. However, attention dispersion is not the only obstacle to effective long-sequence modeling.
@@ 3.3 Representational Preservation and Over-squashing Alleviation
Two other critical challenges in long-context transformers are representational collapse and over-squashing—both fundamentally linked to the properties of softmax attention (Barbero et al., 2024). To clarify these concepts, we provide precise definitions below.
@@
[Representational Collapse and Over-squashing] A transformer model suffers from representational collapse when token representations become similar as sequence length grows: for any $\epsilon>0$, there exists $n_{\epsilon}$ such that $\|\bm{v}_{i}^{(L)}-\bm{v}_{j}^{(L)}\|_{1}<\epsilon$ for all $i,j\leq n$ when $n>n_{\epsilon}$.
@@
Building on these definitions, we show in §C.2 and §C.3 that $\alpha$-entmax addresses both issues: {proposition}[Representational Preservation] There are choices of $\bm{v}^{(0)}\in\mathbb{R}^{n\times d}$ and $\bm{v}^{*(0)}\in\mathbb{R}^{(n+1)\times d}$ such that $\alpha$-entmax ($\alpha>1$) can maintain distinct token representations even as $n\to\infty$: $\|\bm{v}_{n}^{(L)}-\bm{v}_{n+1}^{*(L)}\|_{1}\geq c$ for any constant $c>0$.
@@
[Over-squashing Alleviation] With support size $s\ll n$, $\alpha$-entmax with $\alpha>1$ reduces gradient paths from $\mathcal{O}(n^{L})$ to $\mathcal{O}(s^{L})$, alleviating over-squashing via stronger gradient signals.
To empirically validate these properties, we conducted controlled experiments with transformer architectures employing different attention mechanisms, fully described in §C. For representational collapse, we measured how token representations change when sequence length increases. For over-squashing, we analyzed gradient flow through an 8-layer network on a copying task, where the model must copy information across long distances. The results are shown in Figure 3. The left panel shows the $L_{1}$ norm of representation differences between original and extended sequences after multiple transformer layers. With softmax, this difference rapidly approaches zero as sequence length increases. In contrast, $\alpha$-entmax maintains bounded differences even at extreme lengths (128K tokens), with higher $\alpha$ values providing stronger preservation. The right panel illustrates gradient flow strength across sequence lengths, showing how $\alpha$-entmax maintains significantly stronger gradient signals across layers compared to softmax.
@@ 4.1 Controlling Sparsity in Long Contexts via ASEntmax
Since the logit spread increases only as $\sqrt{\log n}$, a fixed temperature can cause excessive sparsity in long contexts. To address this, we propose Adaptive-Scalable $\alpha$-entmax (ASEntmax), which adaptively scales the temperature for each attention head and query based on the context size and content:
@@
where $\bm{w}_{\beta},\bm{w}_{\gamma}\in\mathbb{R}^{d}$ are learnable, head-specific projection vectors. The scaling factor $(\log n)^{\gamma}$ in Eq. 9 is directly motivated by the asymptotic behavior of normal logits from Remark 4.1. This allows the model learn a slowly rising ($\gamma>0$) or dampening ($\gamma<0$) temperature schedule without interfering in the positional encodings (which would happen with negative values of $\beta$). Specifically, for IID Gaussian logits $\mathcal{N}(0,\sigma)$, when $\delta=0$ and $\gamma=-0.5$ the scaling counteracts the growth of logit ranges ($\Delta_{n}$) as context increases:
@@
which remains constant as $n$ increases, preventing excessive sparsification. Furthermore, with this parameterization, ASEntmax can recover standard $\alpha$-entmax when $\beta=0$, hence allowing a smooth transition between scaled and unscaled regimes. By making $\bm{w}_{\beta}$ and $\bm{w}_{\gamma}$ head-specific and learnable, ASEntmax can adapt to the optimal scaling behavior for each head, balancing the natural concentration benefits of $\alpha$-entmax with precise control over how sparsity patterns evolve with sequence length. Finally, we note that simply scaling the query-key products is appealing from a practical perspective since it allows the direct use of fast optimized kernels for $\alpha$-entmax, such as AdaSplash (Gonçalves et al., 2025), without any modifications.
@@ 4.2 Interacting with Positional Encoding
Positional bias is known to affect attention dynamics and generalization (Barbero et al., 2025; Wu et al., 2025). In this work, we also introduce NAPE (NoPE + ALiBi Positional Encoding), a hybrid scheme in which half of the attention heads receive no positional encoding (NoPE), enabling flexible, data-driven position inference, while the other half use a fixed monotonic bias (ALiBi) to encourage locality. This complementary design allows some heads to focus on global patterns while others encode recency, combining the strengths of both.
While NAPE is not central to our theoretical results, we find it plays a crucial empirical role in enabling robust length generalization, especially when paired with the sparse inductive bias of $\alpha$-attention. As detailed in our theoretical and empirical analysis in §E, the NoPE component of NAPE enables models to develop effective relative positional encoding—supporting the flexibility hypothesis of Kazemnejad et al. (2023)—and, more broadly, $\alpha$-entmax interacts with positional encodings in ways that fundamentally alter attention behavior compared to softmax, such as creating hard attention windows with ALiBi or inducing frequency-dependent patterns with RoPE.
@@ Discussion.
Task-specific patterns further validate our theoretical claims. The consistent superiority of ASEntmax over basic $\alpha$-entmax confirms the benefits of adaptive scaling, particularly at extreme lengths where fixed-$\alpha$ may become too sparse or too diffuse. SSMax performs well on the Copy task, even outperforming other methods, but struggles on more complex tasks like MQMTAR and Reverse at extreme lengths. This indicates that while scaling logits helps maintain peak attention magnitude, the explicit sparsity of $\alpha$-entmax provides additional benefits by completely removing irrelevant connections. These findings are further supported by results on the Max Retrieval task (Figure 5), where sparse attention mechanisms demonstrate superior length extrapolation compared to dense approaches, with ASEntmax maintaining over 60% accuracy even at 4096-length sequences—a dramatic improvement over standard Softmax and Adaptive Temperature (Veličković et al., 2024).
Finally, Copy and Reverse show moderate generalization (up to 64$\times$ and 8$\times$ respectively), while Sort fails beyond 4$\times$ length for all methods. This pattern suggests that tasks requiring precise global ordering (Sort, Reverse) are inherently more challenging for length generalization than tasks dependent on local or independent token properties. We provide per-task results, including results for other positional encoding methods such as ALiBi and NoPE, in §H.
@@ 6 Related Works
The interaction between attention mechanisms and positional encodings significantly impacts long-context generalization. Wu et al. (2025) show that causal masking with distance bias creates a position advantage for early tokens, while Barbero et al. (2025) analyze frequency components in RoPE (Su et al., 2024). Our work shows how $\alpha$-entmax transforms these biases: with NoPE, it helps attention to not always converge to early tokens; with ALiBi, it creates hard attention windows—similar to Hard-ALiBi (Jelassi et al., 2024); with RoPE, it induces frequency-dependent sparsity. Based on this analysis, we introduce NAPE (NoPE + ALiBi), which combines ALiBi’s consistent recency bias with NoPE’s flexible content-based attention for improved long-context generalization. Similar to NAPE, Fu et al. (2025) propose to combine heads with ALiBi and RoPE to increase training stability.
Attention Scaling.
Recent work has shown that scaling attention logits is key for maintaining sharp, focused attention in long contexts. Techniques like YaRN (Peng et al., 2024) and the entropy-aware approach of Zhang et al. (2024) use dynamic logit scaling—often alongside modified RoPE—to stabilize attention when training the model on extend contexts. Scalable-Softmax (Nakanishi, 2025) applies a simple $\log n$ scaling to logits during training to control dispersion without the need for post-training adaptation. InfoScale (Li et al., 2025) derive scaling rules from the principle of entropy invariance, while Scale-invariant Attention (Anson et al., 2025) introduces position-dependent transformations to balance attention across both local and global contexts. Across these methods, adaptive scaling consistently improves extrapolation to longer sequences. Our ASEntmax builds on this line of work by introducing context-dependent, learnable scaling within the $\alpha$-entmax framework, enabling sparse, focused attention as context length increases.
@@
Previous sparse attention approaches include structured patterns like Longformer (Beltagy et al., 2020), BigBird (Zaheer et al., 2020), Flash-Attention (Dao et al., 2022), as well as adaptive methods like $\alpha$-entmax (Peters et al., 2019) and Sparsefinder (Treviso et al., 2022), recently accelerated by AdaSplash (Gonçalves et al., 2025). Our work differs by providing a theoretical analysis of why sparsity helps with long-context generalization, linking it to fundamental limitations including dispersion, representational collapse, and over-squashing.
@@ 7 Conclusions
In this paper, we present a principled approach to long-context modeling by replacing softmax with $\alpha$-entmax in transformer attention. Our theoretical analysis demonstrates how this simple change addresses three fundamental limitations: it avoids attention dispersion through naturally sparse distributions, prevents representational collapse by maintaining distinct token representations, and alleviates over-squashing by reducing gradient paths from $\mathcal{O}(n^{L})$ to $\mathcal{O}(s^{L})$, where $s\ll n$ is the number of tokens with nonzero probability. We further introduce Adaptive-Scalable $\alpha$-entmax (ASEntmax), which adaptively adjusts sparsity based on sequence length for each attention head and query input. Our empirical results confirm these theoretical predictions, with ASEntmax significantly outperforming softmax and existing alternatives across diverse tasks, maintaining high performance even at sequence lengths 1000$\times$ beyond training examples. These findings suggest that addressing the fundamental mathematical limitations of transformer attention mechanisms may provide a direct path to achieve robust long-context generalization.
Limitations.
While our experiments focus on controlled settings with small transformers, the theoretical principles we establish should extend to deeper architectures and larger scales; however, future research on sparse attention in production-scale models is still needed, especially with multi-stage approaches commonly employed for extending the context-length in traditional LLMs (Grattafiori et al., 2024).